Mrs smith earns $1,600 per month and saves $400. what percentage of her monthly income does she save ?
Found 2 solutions by solver91311, jorel555:
Answer by solver91311(24713)



If , then is of .

John

My calculator said it, I believe it, that settles it


Answer by jorel555(1290)

400/1600=1/4=25% of monthly income is saved!!!!!!!!!!